### v1

@@ Abstract
MSC Classification]35R30 (Primary) 68T07, 65N21 (Secondary)
@@ 1 Introduction
Classical methods for solving inverse problems often rely on hand-crafted regularization techniques, explicit models of the forward process, and iterative optimization routines. While mathematically rigorous, these approaches can be limited in their practical application. They tend to be computationally intensive, sensitive to noise, and may struggle to deliver acceptable accuracy when the forward model is not invertible, ill-conditioned, or only partially known. Moreover, designing effective regularization terms requires deep domain expertise and often involves simplifying assumptions that may not hold in realistic scenarios.
In recent times, deep learning has demonstrated an exceptional ability to accurately solve inverse problems and even outperform classical reconstruction methods in linear and nonlinear imaging inverse problems [2, 3, 4]. The subsequent development and success of operator learning methods for the emulation of forward maps of partial differential equations (PDEs) has spurred significant interest in the application of such tools to inverse problems formulated in function spaces [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Indeed, operator learning serves as an attractive alternative and a complement to classical inversion. Architectures such as neural operators are designed to map between infinite-dimensional spaces of functions. As a byproduct, the tunable parameters of these models are decoupled from the resolution of the discretized data used to train them. This property affords exceptional flexibility at inference time and opens up new directions for multiresolution regularization. Operator learning methods can directly approximate the underlying inverse map purely from data. Furthermore, these methods can also be integrated into hybrid approaches that learn regularizers while respecting known physical structure in the forward map.
The chapter distinguishes between solving a single inverse problem with machine learning [19, 20], e.g., requiring the evolution of a Markov chain or running an optimization solver, and solving or probing families of inverse problems simultaneously, which is enabled by operator learning methods. There is an important and popular line of work that employs operator learning to build fast forward map surrogates to be then used downstream in various inverse problems [21, 22, 23, 24, 25, 26, 27, 28]. Under certain conditions, the error incurred by replacing the true forward operator with the surrogate operator controls the propagated error in the inverse problem solution [29, 30, 31, 32, 33]; also see [34, Chp. 3]. However, due to its emphasis on the forward map, such work is outside of the scope of the present chapter. Instead, the chapter focuses on the machine learning of quantities specific to inverse problems and their underlying mathematical structure.
@@
Figure 1: Overall structure of the chapter.
@@ 2 Background on inverse problems
where $\mathcal{G}\colon\mathcal{U}\to\mathcal{Y}$ is a forward operator (also known as the parameter-to-data map), $u\in\mathcal{U}$ is the unknown parameter to be inferred, $y\in\mathcal{Y}$ is the observed data, and $\eta$ models measurement noise or model error. The inverse problem is to recover $u$ from $y$. In many applications arising from science and engineering, $\mathcal{G}$ is nonlinear, the parameter space $\mathcal{U}$ is an infinite-dimensional Banach space, and the data space $\mathcal{Y}$ is high- or infinite-dimensional as well. The inherent nonlinearity and the high- or infinite-dimensional nature of many practical inverse problems can introduce significant mathematical and computational challenges.
More fundamentally, the inverse problem of recovering $u$ from $y$ is often ill-posed in the sense of Hadamard: solutions may not exist, may not be unique, or may not depend continuously on the data [1]. Regarding the last difficulty, the inverse problem solution’s sensitivity to noise in $y$ makes naïve inversion schemes unstable and unreliable. To address this challenge, one may introduce additional information, such as smoothness, sparsity, or statistical structure, through regularization techniques or by adopting a probabilistic framework to capture uncertainty and stabilize the solution. This section reviews approaches in both directions.
A canonical example of a severely ill-posed nonlinear inverse problem is electrical impedance tomography (EIT) [43]. EIT is a noninvasive imaging technique used to reconstruct the internal conductivity field of an object by making electrical measurements on the object’s boundary. Spatial variations in the conductivity correspond to different regions of interest in the object, such as malignant biological tissue in hospital patients or distinct Earth subsurface structures in a geophysical context. Due to its well-understood mathematical properties, EIT has emerged as a prototypical nonlinear benchmark inverse problem for both traditional inverse solvers and deep learning methods [44, 45, 46, 47, 48, 13, 12, 45, 49, 50, 51, 52]. This chapter works with the idealized continuum measurement version of the problem, which is known as Calderón’s problem [53]. For convenience, we will use the Calderón problem and EIT terminology interchangeably. The following EIT example will appear frequently throughout the chapter.
(a) $\{(g_{n},\Lambda_{\gamma}g_{n})\}_{n=1}^{2}$
@@
Figure 2: Electrical impedance tomography from Example 1. The domain $\Omega$ is the unit disk in $\mathbb{R}^{2}$. The boundary manifold $\partial\Omega$ is identified as the one-dimensional unit torus $\mathbb{T}$ by transforming into polar coordinates. Figure 2(a) displays two pairs of randomly sampled Neumann data ($g_{n}$, black) and corresponding Dirichlet data ($\Lambda_{\gamma}g_{n}$, red) on the torus. In Fig. 2(b), the full NtD map $\Lambda_{\gamma}$ is represented numerically using its $L^{2}(\partial\Omega\times\partial\Omega)$ integral kernel function; this kernel function is further shifted, re-scaled, and transformed to the torus to aid visualization. The true realization of the conductivity $\gamma$ generating these measurements under the forward map $\gamma\mapsto\Lambda_{\gamma}$ is shown in Fig. 2(c). Solving the inverse conductivity problem amounts to evaluating the inverse map $\Lambda_{\gamma}\mapsto\gamma$.
@@
EIT aims to recover an unknown conductivity field $u=\gamma\in L^{\infty}(\Omega;\mathbb{R}_{>0})$ from measurements corresponding to the Neumann-to-Dirichlet (NtD) map $\mathcal{G}(\gamma)\coloneqq\Lambda_{\gamma}$ of the elliptic PDE
@@
The NtD map $\Lambda_{\gamma}\in\mathcal{L}(H^{-1/2}(\partial\Omega);H^{1/2}(\partial\Omega))$ is a continuous linear operator from the dual Sobolev space $H^{-1/2}(\partial\Omega)=(H^{1/2}(\partial\Omega))^{*}$ into the trace Sobolev space $H^{1/2}(\partial\Omega)$ of $H^{1}(\Omega)$ boundary values [54]. It maps the current density $g$ to the voltage $\phi|_{\partial\Omega}$, i.e., $\Lambda_{\gamma}g=\phi|_{\partial\Omega}$ according to (2.2), both of which are functions on the smooth boundary $\partial\Omega$ of a bounded domain $\Omega\subset\mathbb{R}^{d}$. Figure 2 illustrates the main components of the EIT problem.
The problem of finding $\gamma$ from $\Lambda_{\gamma}$ is hard. Indeed, we wish to recover a function, often in an infinite-dimensional space, on the interior of $\Omega$ only by probing the boundary $\partial\Omega$. More quantitatively, EIT is severely ill-posed from a stability perspective. That is, for sufficiently smooth conductivities $\gamma_{0}$ and $\gamma_{1}$ bounded away from zero and infinity, it holds that
@@
whenever $\delta=\lVert\Lambda_{\gamma_{0}}-\Lambda_{\gamma_{1}}\rVert_{\mathcal{L}(H^{-1/2}(\partial\Omega);H^{1/2}(\partial\Omega))}$ is small enough [55]. This means that for the parameter error $\lVert\gamma_{0}-\gamma_{1}\rVert_{L^{\infty}(\Omega)}$ to be less than $\varepsilon$, it suffices for the measurement error $\delta$ to be exponentially small: $\delta\lesssim\exp(-\varepsilon^{-1/\beta})$. This estimate only holds in the range of the forward map $\mathcal{G}\colon\gamma\mapsto\Lambda_{\gamma}$. See Fig. 3 for a visualization. Similar results remain valid outside of the range of $\mathcal{G}$, as will be discussed in Sections 4.1.2 and 4.3.1. These extensions are useful to handle realistic noisy measurements.
@@ 2.1.1 Bayesian inverse problems
The Bayesian approach to inverse problems is perhaps the most well-developed of all the probabilistic methods mentioned earlier. In recent decades, there has been significant progress and success in integrating Bayesian inference with statistical inverse problems formulated in function spaces [57, 58]. This has led to comprehensive well-posedness theory, approximation theory, and posterior consistency theory [59, 60] on the one hand, and scalable computational algorithms on the other hand [61, 62]. By leveraging an inherently infinite-dimensional framing, the Bayesian approach is naturally applicable to inverse problems from continuum applied mathematics and scientific computing.
The Bayesian statistician models all quantities in the inverse problem (2.1) probabilistically. The noise $\eta$ is assumed to be a random variable with distribution $\nu_{0}\coloneqq\operatorname{Law}(\eta)$. This and the forward map $\mathcal{G}$ defines the likelihood of observing $y$ given the parameter $u$; we denote by $\nu_{\mathcal{G}(u)}\coloneqq\operatorname{Law}(y\,|\,u)$ the pushforward of $\nu_{0}$ under the translation-by-$\mathcal{G}(u)$ map. The core of any Bayesian procedure is the prior distribution specification. The parameter $u$ is modeled as a random variable with distribution $\mu\in\mathscr{P}(\mathcal{U})$, the prior. The prior probability measure $\mu$ represents the modeler’s beliefs and existing knowledge before observing any data. The prior $\mu$ is known to regularize the ill-posedness of the inverse problem. Popular choices in infinite dimensions include Gaussian priors, Besov priors, or data-driven pushforwards thereof [60].
The Bayesian solution to the inverse problem is not just a single point estimate of the parameter $u\in\mathcal{U}$, but instead the entire posterior probability distribution $\mu^{y}\coloneqq\operatorname{Law}(u\,|\,y)\in\mathscr{P}(\mathcal{U})$ obtained by conditioning the prior on the random variable $y$ according to (2.1). Assume that $\nu_{\mathcal{G}(u)}$ has density $\exp(-\Phi(u;{\,\cdot\,}))$ with respect to $\nu_{0}$ for all $u$, $\mu$ almost everywhere (a.e.). Then, under some mild conditions that can be verified on a case-by-case basis, the infinite-dimensional Bayes’ rule [60, Thm. 14, pp. 9–10] shows that the posterior is given by
@@ 2.1.2 Transport maps
Computational measure transport is a general paradigm that seeks to represent a target probability measure $\nu=\mathsf{T}^{\star}_{\#}\rho$ as the pushforward of some easy-to-sample reference probability measure $\rho$ under a measurable function $\mathsf{T}^{\star}$ [79]. Thus, if $u\sim\rho$ is a sample from $\rho$, then $y=\mathsf{T}^{\star}(u)\sim\nu$ is a sample from $\nu$. Such a task can be accomplished by solving the minimum divergence estimation problem
@@
Specializing to the Bayesian setting discussed in Section 2.1.1, to solve inverse problems with measure transport, one typically seeks a prior-to-posterior map $\mathsf{T}=\mathsf{T}^{(\mu,y)}$ such that $\mu^{y}=\mathsf{T}_{\#}\mu$ [34, Chp. 5], where $\mu^{y}$ is the posterior (2.4) corresponding to prior $\mu$ and observed data $y$. The dependence of $\mathsf{T}$ on $\mu$ and $y$ is implicit from the previous equality. While not usually done, in some applications it makes sense to explicitly parametrize the dependence of $\mathsf{T}$ on $\mu$ or $y$. This way, problem (2.5) does not have to be re-solved every time $\mu$ or $y$ is changed, which can be prohibitively expensive. There is work along these lines that develops amortized solvers for Bayesian inverse problems [82, 81]. This amounts to methods that, given a new piece of data $y$, can rapidly produce a representation of the posterior distribution corresponding to that $y$ without re-training or additional optimization. Using measure transport, this is equivalent to approximating the map $y\mapsto\mathsf{T}^{(y)}$, where $\mathsf{T}^{(y)}$ is the divergence minimization solution for a particular $y$. The amortized estimation of the data-to-posterior map is arguably a more challenging task than merely emulating the data-to-parameter inverse map, because now the output space is the metric space $\mathscr{P}(\mathcal{U})$ of probability measures supported on $\mathcal{U}$, rather than just the linear parameter space $\mathcal{U}$ itself. We return to data-to-posterior estimation in Section 4.4.

### v2

@@ Abstract
MSC Classification]35R30 (Primary); 68T07, 65N21 (Secondary)
@@ 1 Introduction
Classical methods for solving inverse problems often rely on hand-crafted regularization techniques, explicit models of the forward process, and iterative optimization routines. These methods are mathematically rigorous, well-established, and offer interpretability and theoretical guarantees. In practice, they may require substantial computation and careful modeling. Their performance can be sensitive to noise or to forward models that are noninvertible, ill-conditioned, or only partially specified.
In recent times, deep learning has demonstrated an exceptional ability to accurately solve inverse problems and even outperform classical reconstruction methods in linear and nonlinear imaging inverse problems [2, 3, 4]. The subsequent development and success of operator learning methods for the emulation of forward maps of partial differential equations (PDEs) has spurred significant interest in the application of such tools to inverse problems formulated in function spaces [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Indeed, operator learning serves as an attractive alternative and a complement to classical inversion. Architectures such as neural operators are designed to map between infinite-dimensional spaces of functions. As a byproduct, the tunable parameters of these models are decoupled from the resolution of the discretized data used to train them. This property affords exceptional flexibility at inference time and opens up new directions for multiresolution regularization. This chapter focuses on applying operator learning methods through two complementary strategies: (i) directly approximating the underlying inverse map from data alone, referred to as “end-to-end learning,” and (ii) learning regularizers that respect known physical structure in the forward map, to then be incorporated into existing inverse solvers requiring regularization.
The chapter distinguishes between solving a single inverse problem with machine learning [19, 20, 21, 22], e.g., requiring the evolution of a Markov chain or running an optimization solver, and simultaneously solving or probing families of inverse problems with the same forward model but different observed data, which is enabled by operator learning methods. There is an important and popular line of work that employs operator learning to build fast forward map surrogates for subsequent downstream use in various inverse problems [23, 24, 25, 26, 27, 28, 29, 30]. Under certain conditions, the error incurred by replacing the true forward operator with the surrogate operator controls the propagated error in the inverse problem solution [31, 32, 33, 34, 35]; also see [36, Chp. 3]. However, due to its emphasis on the forward map, such work is outside of the scope of the present chapter. Instead, the chapter focuses on the machine learning of quantities specific to inverse problems and their underlying mathematical structure.
@@
Figure 1: Overall structure of the chapter. The dotted lines indicate the connections associated with end-to-end learning, whereas the dashed lines correspond to regularizer learning.
@@ 2 Background on inverse problems
where $\mathcal{G}\colon\mathcal{U}\to\mathcal{Y}$ is a forward operator (also known as the parameter-to-data map), $u\in\mathcal{U}$ is the unknown parameter to be inferred, $y\in\mathcal{Y}$ is the observed data, and $\eta$ models measurement noise or model error. The inverse problem is to recover $u$ from $y$. In many applications arising from science and engineering, $\mathcal{G}$ is nonlinear. The parameter space $\mathcal{U}$ is contained in an infinite-dimensional Banach space; this is referred to as a nonparametric setting. The data space $\mathcal{Y}$ is high- or infinite-dimensional as well. The inherent nonlinearity and the high- or infinite-dimensional nature of many practical inverse problems can introduce significant mathematical and computational challenges.
More fundamentally, the inverse problem of recovering $u$ from $y$ is often ill-posed in the sense of Hadamard: solutions may not exist, may not be unique, or may not depend continuously on the data [1]. Regarding the last difficulty, the inverse problem solution’s sensitivity to noise in $y$ makes naïve inversion schemes unstable and unreliable. To address this challenge, one may introduce additional information about the parameter $u$ in (2.1), such as smoothness, sparsity, or statistical structure. This can be done either through regularization techniques or by adopting a probabilistic framework to capture uncertainty and stabilize the solution. This section reviews approaches in both directions.
A canonical example of a severely ill-posed nonlinear inverse problem is electrical impedance tomography (EIT) [45]. EIT is a noninvasive imaging technique used to reconstruct the internal conductivity field of an object by making electrical measurements on the object’s boundary. Spatial variations in the conductivity correspond to different regions of interest in the object, such as malignant biological tissue in hospital patients or distinct Earth subsurface structures in a geophysical context. Due to its well-understood mathematical properties, EIT has emerged as a prototypical nonlinear benchmark inverse problem for both traditional inverse solvers and deep learning methods [46, 47, 48, 49, 13, 12, 50, 51, 52, 53, 54]. This chapter works with the idealized continuum measurement version of the problem, which is known as Calderón’s problem [55]. For convenience, we will use the Calderón problem and EIT terminology interchangeably. The following EIT example will appear frequently throughout the chapter.
(a) $\{(g_{m},\Lambda_{\gamma}g_{m})\}_{m=1}^{2}$
@@
Figure 2: Electrical impedance tomography from Example 1. The domain $\Omega$ is the unit disk in $\mathbb{R}^{2}$. The boundary manifold $\partial\Omega$ is identified as the one-dimensional unit torus $\mathbb{T}$ by transforming into polar coordinates. Figure 2(a) displays two pairs of randomly sampled Neumann data ($g_{m}$, black) and corresponding Dirichlet data ($\Lambda_{\gamma}g_{m}$, red) on the torus. In Fig. 2(b), the full NtD map $\Lambda_{\gamma}$ is represented numerically using its $L^{2}(\partial\Omega\times\partial\Omega)$ integral kernel function; this kernel function is further shifted, re-scaled, and transformed to $\mathbb{T}^{2}$ to aid visualization. The true realization of the conductivity $\gamma$ generating these measurements under the forward map $\gamma\mapsto\Lambda_{\gamma}$ is shown in Fig. 2(c). Solving the inverse conductivity problem amounts to evaluating the inverse map $\Lambda_{\gamma}\mapsto\gamma$.
@@
The mathematical model of EIT begins with the linear elliptic PDE
@@
In (2.2), $\Omega\subset\mathbb{R}^{d}$ is the medium of interest, $\gamma\in L^{\infty}(\Omega)$ is the strictly positive almost everywhere electrical conductivity field, $\phi$ is the electric potential, $g$ is the applied boundary current, and $\partial/\partial\mathsf{n}$ is the outward normal derivative to the boundary $\partial\Omega$. We assume that $\partial\Omega$ is smooth. The PDE defines a solution operator that maps a given pair $(\gamma,g)$ to $\phi$ solving (2.2). However, in EIT, interior quantities such as $\phi$ are not accessible. Instead, for a fixed $\gamma$, one can vary the boundary current density $g$ and measure the corresponding boundary voltages $\phi|_{\partial\Omega}$. This measurement process leads to a linear operator $\Lambda_{\gamma}$ called the Neumann-to-Dirichlet (NtD) map, which is defined by
$g\mapsto\Lambda_{\gamma}g\coloneqq\phi|_{\partial\Omega}$ (2.3)
subject to (2.2). It is known that the NtD map $\Lambda_{\gamma}\in\mathcal{L}(H^{-1/2}(\partial\Omega);H^{1/2}(\partial\Omega))$ is a continuous linear operator from the dual Sobolev space $H^{-1/2}(\partial\Omega)=(H^{1/2}(\partial\Omega))^{*}$ into the trace Sobolev space $H^{1/2}(\partial\Omega)$ of $H^{1}(\Omega)$ boundary values [56]. In the notation of (2.1), we take $\mathcal{U}\coloneqq L^{\infty}(\Omega)$, $\mathcal{Y}\coloneqq\mathcal{L}(H^{-1/2}(\partial\Omega);H^{1/2}(\partial\Omega))$, $u\coloneqq\gamma$, and $\mathcal{G}(\gamma)\coloneqq\Lambda_{\gamma}$. The Calderón inverse problem is to reconstruct the unknown $\gamma$ from $\Lambda_{\gamma}$. Figure 2 illustrates the main components of EIT.
The inverse problem of finding $\gamma$ from $\Lambda_{\gamma}$ is hard. Indeed, we wish to recover a function, often in an infinite-dimensional space, on the interior of $\Omega$ only by probing the boundary $\partial\Omega$. More quantitatively, EIT is severely ill-posed from a stability perspective. That is, for sufficiently smooth conductivities $\gamma_{0}$ and $\gamma_{1}$ bounded away from zero and infinity, it holds for some $\beta>0$ that
@@
whenever $\delta=\lVert\Lambda_{\gamma_{0}}-\Lambda_{\gamma_{1}}\rVert_{\mathcal{L}(H^{-1/2}(\partial\Omega);H^{1/2}(\partial\Omega))}$ is small enough [57, 58]. This means that for the parameter error $\lVert\gamma_{0}-\gamma_{1}\rVert_{L^{\infty}(\Omega)}$ to be less than $\varepsilon$, it suffices for the measurement error $\delta$ to be exponentially small: $\delta\lesssim\exp(-\varepsilon^{-1/\beta})$. This estimate only holds in the range of the forward map $\mathcal{G}\colon\gamma\mapsto\Lambda_{\gamma}$. See Fig. 3 for a visualization. Results similar to (2.4) remain valid outside of the range of $\mathcal{G}$, as Sections 4.1.2 and 4.3.1 will discuss. These extensions are useful to handle realistic noisy measurements [59].
@@ 2.1.1 Bayesian inverse problems
The Bayesian approach to inverse problems is perhaps the most well-developed of all the probabilistic methods mentioned earlier. In recent decades, there has been significant progress and success in integrating Bayesian inference with statistical inverse problems formulated in function spaces [61, 62]. This has led to comprehensive well-posedness theory, approximation theory, and posterior consistency theory [63, 64] on the one hand, and scalable computational algorithms on the other hand [65, 66]. By leveraging an inherently infinite-dimensional framing, the nonparametric Bayesian approach is naturally applicable to inverse problems from continuum applied mathematics and scientific computing.
The Bayesian statistician models all quantities in the inverse problem (2.1) probabilistically. The noise $\eta$ is assumed to be a random variable with distribution $\nu_{0}\coloneqq\operatorname{Law}(\eta)$. This and the forward map $\mathcal{G}$ define the likelihood of observing $y$ given the parameter $u$; we denote by $\nu_{\mathcal{G}(u)}\coloneqq\operatorname{Law}(y\,|\,u)$ the pushforward of $\nu_{0}$ under the translation-by-$\mathcal{G}(u)$ map. The core of any Bayesian procedure is the prior distribution specification. The parameter $u$ is modeled as a random variable with distribution $\mu\in\mathscr{P}(\mathcal{U})$, the prior. The prior probability measure $\mu$ represents the modeler’s beliefs and existing knowledge before observing any data. The prior $\mu$ is known to regularize the ill-posedness of the inverse problem. Popular choices in infinite dimensions include Gaussian priors, Besov priors, or data-driven pushforwards thereof [64].
The Bayesian solution to the inverse problem is not just a single point estimate of the parameter $u\in\mathcal{U}$, but instead the entire posterior probability distribution $\mu^{y}\coloneqq\operatorname{Law}(u\,|\,y)\in\mathscr{P}(\mathcal{U})$ obtained by conditioning the prior on the random variable $y$ according to (2.1). Assume that $\nu_{\mathcal{G}(u)}$ has density $\exp(-\Phi(u;{\,\cdot\,}))$ with respect to $\nu_{0}$ for all $u$, $\mu$ almost everywhere (a.e.). Then, under some mild conditions that can be verified on a case-by-case basis, the infinite-dimensional Bayes’ rule [64, Thm. 14, pp. 9–10] shows that the posterior is given by
@@ 2.1.2 Transport maps
Computational measure transport is a general paradigm that seeks to represent a target probability measure $\nu=\mathsf{T}^{\star}_{\#}\rho$ as the pushforward of some easy-to-sample reference probability measure $\rho$ under a measurable function $\mathsf{T}^{\star}$ [83]. Thus, if $u\sim\rho$ is a sample from $\rho$, then $y=\mathsf{T}^{\star}(u)\sim\nu$ is a sample from $\nu$. Such a task can be approximately accomplished by solving the minimum divergence estimation problem
@@
Specializing to the Bayesian setting discussed in Section 2.1.1, to solve inverse problems with measure transport, one typically seeks a prior-to-posterior map $\mathsf{T}=\mathsf{T}^{(\mu,y)}$ such that $\mu^{y}=\mathsf{T}_{\#}\mu$ [36, Chp. 5], where $\mu^{y}$ is the posterior (2.5) corresponding to prior $\mu$ and observed data $y$. The dependence of $\mathsf{T}$ on $\mu$ and $y$ is implicit from the previous equality. While not usually done, in some applications it makes sense to explicitly parametrize the dependence of $\mathsf{T}$ on $\mu$ or $y$. This way, problem (2.6) does not have to be re-solved every time $\mu$ or $y$ is changed, which can be prohibitively expensive. There is work along these lines that develops amortized solvers for Bayesian inverse problems [86, 85]. This amounts to methods that, given a new piece of data $y$, can rapidly produce a representation of the posterior distribution corresponding to that $y$ without re-training or additional optimization. Using measure transport, this is equivalent to approximating the map $y\mapsto\mathsf{T}^{(y)}$, where $\mathsf{T}^{(y)}$ is the divergence minimization solution for a particular $y$. The amortized estimation of the data-to-posterior map is arguably a more challenging task than merely emulating the data-to-parameter inverse map, because now the output space is the metric space $\mathscr{P}(\mathcal{U})$ of probability measures supported on $\mathcal{U}$, rather than just the linear parameter space $\mathcal{U}$ itself. We return to data-to-posterior estimation in Section 4.4.